# Machine Learning, ICML

### v1

@@ Abstract
Estimating test accuracy without access to the ground-truth test labels under varying test environments is a challenging, yet extremely important problem in the safe deployment of machine learning algorithms. Existing works rely on the information from either the outputs or the extracted features of neural networks to formulate an estimation score correlating with the ground-truth test accuracy. In this paper, we investigate–both empirically and theoretically–how the information provided by the gradients can be predictive of the ground-truth test accuracy even under a distribution shift. Specifically, we use the norm of classification-layer gradients, backpropagated from the cross-entropy loss after only one gradient step over test data. Our key idea is that the model should be adjusted with a higher magnitude of gradients when it does not generalize to the test dataset with a distribution shift. We provide theoretical insights highlighting the main ingredients of such an approach ensuring its empirical success. Extensive experiments conducted on diverse distribution shifts and model structures demonstrate that our method significantly outperforms state-of-the-art algorithms.
Machine Learning, ICML
@@ 1 Introduction
Deploying machine learning models in the real world is often subject to a distribution shift between training and test data. Such a shift may significantly degrade the model’s performance during test time (Quinonero-Candela et al., 2008; Geirhos et al., 2018; Koh et al., 2021) and lead to high risks related to AI satefy (Deng & Zheng, 2021). To alleviate this problem, a common practice is to monitor the model performance regularly via collecting the ground truth of a subset of the current test dataset (Lu et al., 2023). This practice, however, is usually time-consuming and expensive highlighting the importance of proposing unsupervised methods for estimating test performance under distribution shift, commonly known as unsupervised accuracy estimation.
@@
Current studies mainly focus on outputs or feature representation to derive an test error estimation score. Such a score represents calibrated test error or distribution discrepancy from the training to the test datasets (Hendrycks & Gimpel, 2016; Guillory et al., 2021; Garg et al., 2022; Deng & Zheng, 2021; Yu et al., 2022b; Lu et al., 2023). For example, Hendrycks & Gimpel (2016) calculates the average maximum softmax score of the test samples as the estimated error. Similarly, Garg et al. (2022) learns a confidence threshold from the training distribution. Deng & Zheng (2021) quantifies the distribution difference between the training and the test datasets in the feature space, while Yu et al. (2022b) gauges the distribution gap in the parameter level. Although insightful, the current body of literature overlooks another important aspect of test accuracy estimation even though gradients are known to correlate strongly with the generalization error of the deep neural networks (Li et al., 2019; An et al., 2020). Thus, in this work, we seek to answer the following question:
Are gradients predictive of test errors under distribution shift?
@@
The role of gradients in generalization has attracted increasing attention, and many algorithms are designed based on them (Shi et al., 2021; Zhou et al., 2020; Mansilla et al., 2021; Zhao et al., 2022; Guiroy et al., 2019). For example, Mansilla et al. (2021) clips the conflicting gradients emerging in domain shift scenarios and promote gradient agreement across multiple domains via the gradient agreement strategy. Meanwhile, Zhao et al. (2022) designs a gradient-based regularization term to make the optimizer find the flat minima. In a more general field, Huang et al. (2021) designs a gradient-based function estimator to detect out-of-distribution samples (comparison is shown in Appendix F). However, despite their empirical success, how gradients affect generalization performance is still an open question. In this paper, we surprisingly observe that there exists a strong linear relationship between the gradient norm and the test accuracy under distribution shift. To explore this phenomenon, we theoretically certify that well-calibrated gradients include information directly related to the test accuracy. Overall, this work provides a direct evidence for the correlation between gradient-based information and generalization performance, which might help us identify how and when neural networks make errors as well as generalizing across unseen domains.
@@
We tackle the above question by transposing the idea of using the gradient norm for generalization accuracy prediction to the distribution-shift setting. We hypothesize that the model requires a large magnitude gradient step if it cannot generalize well on the test dataset from unseen domains at hand. To quantify the magnitude of gradients, we propose a simple yet efficient gradient-based statistic, GdScore, which employs the vector norm of gradients backpropagated from a standard cross-entropy loss on the test samples. To avoid the need for ground-truth labels, we propose a pseudo-labeling strategy that seeks to benefit from both correct and incorrect predictions. Ultimately, we demonstrate that the simple one-step gradient norm of the classification layer strongly correlates with the generalization performance under diverse distribution shifts acting as a strong and lightweight proxy for the latter. The main contributions of our paper are summarized as follows:
@@
    We first provide several theoretical insights showing that correct pseudo-labeling and gradient norm have a direct impact on the test error estimation. This is achieved by analyzing the analytical expression of the gradient after 1 gradient step over the trained model on test data under distribution shift and by upper-bounding the population out-of-distribution risk.
@@
    Based on these theoretical insights, we propose the GdScore, which gauges the magnitude of the classification-layer gradients and shows a strong correlation with test accuracy. Our method does not require access to neither test labels nor training datasets, and only needs one step of backpropagation making it particularly lightweight in terms of computational efficiency compared with other self-training methods.
@@
    We perform a large-scale empirical evaluation of GdScore showing its superior performance. We achieve new state-of-the-art results on 11 benchmarks across diverse distribution shifts compared to 8 competitors, while being faster than the previous best baseline.
The rest of the paper is organized as follows. Section 2 presents the necessary background on the problem setup of our work. In Section 3, we derive several theoretical insights that motivate the GdScore introduced afterward. Section 4 is devoted to extensive empirical evaluation of our method, while ablation study is deferred to Section 5. Finally, Section 6 concludes our work.
2 Background knowledge
@@
where $\mathrm{s}_{\bm{\omega}}^{(k)}$ denotes the softmax function for class $k$ that outputs an approximation of the posterior probability $P(Y\!=\!k|\mathbf{x})$: $\mathrm{s}_{\bm{\omega}}^{(k)}(f_{\mathbf{g}}(\mathbf{x}))=\exp\{\mathbf{w}_{k}^{\intercal}f_{\mathbf{g}}(\mathbf{x})\}/\left(\sum_{\tilde{k}}\exp\{\mathbf{w}_{\tilde{k}}^{\intercal}f_{\mathbf{g}}(\mathbf{x})\}\right)$.
@@
We now assume to have access to $m$ test samples from the target distribution $\mathcal{D}_{\text{test}}=\{\tilde{\mathbf{x}_{i}}\}_{i=1}^{m}\sim P_{T}(\mathbf{x})$, where $P_{T}(\mathbf{x},y)\neq P_{S}(\mathbf{x},y)$. For each test sample $\tilde{\mathbf{x}}_{i}$, we predict the label by $\tilde{y}^{\prime}_{i}=\operatorname*{arg\,max}_{k\in\mathcal{Y}}f_{\bm{\theta}}(\tilde{\mathbf{x}}_{i})$. We now want to assess the performance of $f_{\bm{\theta}}$ on a target distribution without using corresponding ground-truth labels $\{\tilde{y}\}_{i=1}^{m}$ by estimating as accurately as possible the following quantity:
@@
3 GdScore: on the strong correlation between gradient norm and test accuracy
We start by deriving an analytical expression of the gradient obtained when fine-tuning a source pre-trained model on new test data. We further use the intuition derived from it to propose our test accuracy estimation score, and justify its effectiveness through a more thorough theoretical analysis.
@@ 3 GdScore: on the strong correlation between gradient norm and test accuracy
Below, we follow the setup considered by Denevi et al. (2019); Balcan et al. (2019); Arnold et al. (2021) to develop our intuition behind the importance of gradient norm in unsupervised accuracy estimation. Our main departure point for this analysis is to consider fine-tuning: a popular approach to adapting a pre-trained model to new labeled datasets is to update either all or just a fraction of its parameters using gradient descent on the new data. To this end, let us consider the following linear regression example, where the test data from unseen domains are distributed as $X\!\sim\!\mathcal{N}(0,\sigma^{x}_{t})$, $(Y|X\!=\!x)\!\sim\!\mathcal{N}(\theta_{t}x,1)$ parameterized by the optimal regressor $\theta_{t}\in\mathbb{R}$, while the data on which the model was trained is distributed as $X\!\sim\!\mathcal{N}(0,\sigma^{x}_{s})$, $(Y|X\!=\!x)\!\sim\!\mathcal{N}(\theta_{s}x,1)$ with $\theta_{s}\in\mathbb{R}$. Consider the least-square loss over the test distribution:
@@
$\quad\frac{1}{2}\nabla_{c}\mathbb{E}_{P_{T}(x)}\mathbb{E}_{P_{S}(y|x)}[(y-cx)^{2}]$
$=\mathbb{E}_{P_{T}(x)}\mathbb{E}_{P_{S}(y|x)}[(y-cx)(-x)]$
@@
3.1 Proposed approach
We now formally introduce our proposed score, termed GdScore, to estimate test accuracy in an unsupervised manner during evaluation. We start by recalling the backpropagation process of the pre-trained neural network $f_{\theta}$ from a cross-entropy loss and then describe how to leverage the gradient norm for the unsupervised accuracy estimation. The detailed algorithm can be found in Appendix A.
@@ 3.1 Proposed approach
where $\tau$ denotes the threshold value, and $U[1,k]$ denotes the discrete uniform distribution with outcomes $\{1,\dots,K\}$. In a nutshell, we assign the predicted label to $\tilde{\mathbf{x}}_{i}$, when the prediction confidence is larger than a threshold while using a randomly sampled label from the label space otherwise. The detailed empirical evidence justifying this choice is shown in Section 5, and we discuss the choice of proper threshold $\tau$ in Appendix B. From the theoretical point of view, our approach assumes that the classifier makes mistakes mostly on data with low prediction confidence, for which we deliberately assign noisy pseudo-labels. Feofanov et al. (2019) used a similar approach to derive an upper bound on the test error and proved its tightness in the case where the assumption is satisfied. We discuss this matter in more detail in Appendix G.
@@
To estimate our score, we calculate the gradients w.r.t. the weights of the classification layer during the first epoch backpropagated over the standard cross-entropy loss. The loss function can be written as follows:
@@
Given parameters of the classification layer $\bm{\omega}$, the gradients of the standard cross-entropy loss are:
@@
Note that our method requires neither gradients of the whole parameter set of the pre-trained model nor iterative training hinting at its high computational efficiency.
@@
where $||\cdot||_{p}$ denotes $L_{p}$-norm, and $\bm{\omega}$ is the classification component of the network of parameters $\bm{\theta}$.
@@ 3.2 Theoretical analysis
In this section, we provide some theoretical insights into our method. We first clarify the connection between the true target cross-entropy error and the norm of the gradients. Then, we show that the gradient norm is upper-bounded by a weighted sum of the norm of the inputs.
@@
where $\bm{\omega}=(\mathbf{w}_{k})_{k=1}^{K}\in\mathbb{R}^{D\times K}$ are the parameters of the linear classification layer $f_{\bm{\omega}}$. For the ease of notation, the gradient of $\mathcal{L}_{T}$ w.r.t $\bm{\omega}$ is denoted by $\nabla\mathcal{L}_{T}$. The following theorem, whose proof we defer to Appendix I.1, makes the connection between the true risk and the $L_{p}$-norm of the gradient explicit.
@@
The left-hand side here is the difference in terms of the true risks obtained for the same distribution with two different classifiers. The right-hand side shows how this difference is controlled by the maximum gradient norm over the two classifiers and a term capturing how far the two are apart. In the context of the proposed approach, we want to know the true risk of the source classifier $\bm{\omega}_{s}$ on the test data and its change after one step of gradient descent. The following corollary applies Theorem 3.1 to characterize this exact case. The proof is deferred in Appendix I.2.
@@
$|\mathcal{L}_{T}(\bm{\omega}_{s})-\mathcal{L}_{T}(\mathbf{c})|$ $\leq\eta\cdot\max_{\bm{\omega}}(\left\lVert\nabla\mathcal{L}_{T}(\bm{\omega})\right\rVert_{p})\cdot\left\lVert\nabla\mathcal{L}_{T}(\bm{\omega}_{s})\right\rVert_{q}.$
Note that in this case $\mathcal{L}_{T}(\bm{\omega}_{s})$ can be seen as a term providing the true test risk after pseudo-labeling it with the source classifier. This shows the importance of pseudo-labeling as it acts as a departure point for obtaining a meaningful estimate of the right-hand side. When the latter is meaningful, the gradient norm on the right-hand side controls it together with a magnitude that tells us how far we went after one step of backpropagation. In the next theorem, we provide an upper bound on the $L_{p}$-norm of the gradient as a weighted sum of the $L_{p}$-norm of the inputs. The proof is deferred in Appendix I.3.
@@
For any $p\geq 1$, the $L_{p}$-norm of the gradient can be upper-bounded as follows:
$\left\lVert\nabla\mathcal{L}_{T}(\bm{\omega})\right\rVert_{p}$ $\leq\mathbb{E}_{P_{T}(\mathbf{x},y)}\alpha(\bm{\omega},\mathbf{x},y)\cdot\left\lVert\mathbf{x}\right\rVert_{p},\quad\forall\bm{\omega}\in\mathbb{R}^{D\times K}$
@@
Hence, the norm of the gradient is upper-bounded by a weighted combination of the norm of the inputs, where the weight $\alpha(\bm{\omega},\mathbf{x},y)\in[0,1]$ conveys how well the model predicts on $\mathbf{x}$. In the case of perfect classification, the upper bound is tight and equal to 00. In practice, as we do not have access to the true risk, its gradient can be approximated by the proposed GdScore and pseudo-label test data by Eq. 5. As we said earlier, this requires the model to be well calibrated, which is conventional to assume for self-training methods (Amini et al., 2023) including the approach of Yu et al. (2022b). Then, the network projects test data into the low confidence regions, and the gradient for these examples will be large as we need to update $\bm{\omega}_{s}$ significantly to fit them.

### v3

# Leveraging Gradients for Unsupervised Accuracy Estimation under Distribution Shift

@@ Abstract
Estimating the test performance of a model, possibly under distribution shift, without having access to the ground-truth labels is a challenging, yet very important problem for the safe deployment of machine learning algorithms in the wild. Existing works mostly rely on information from either the outputs or the extracted features of neural networks to estimate a score that correlates with the ground-truth test accuracy. In this paper, we investigate – both empirically and theoretically – how the information provided by the gradients can be predictive of the ground-truth test accuracy even under distribution shifts. More specifically, we use the norm of classification-layer gradients, backpropagated from the cross-entropy loss after only one gradient step over test data. Our intuition is that these gradients should be of higher magnitude when the model generalizes poorly. We provide the theoretical insights behind our approach and the key ingredients that ensure its empirical success. Extensive experiments conducted with various architectures on diverse distribution shifts demonstrate that our method significantly outperforms current state-of-the-art approaches. The code is available at https://github.com/Renchunzi-Xie/GdScore.
@@ 1 Introduction
Deploying machine learning models in the real world is often subject to a distribution shift between training and test data. Such a shift may significantly degrade the model’s performance at test time (Quinonero-Candela et al., 2008; Geirhos et al., 2018; Koh et al., 2021) and lead to high risks related to AI satefy (Deng & Zheng, 2021). To alleviate this problem, a common practice is to monitor the model performance regularly by collecting the ground truth of a subset of the current test dataset (Lu et al., 2023). However, this is usually time-consuming and expensive, highlighting the need for unsupervised methods to assess the test performance of models under distribution shift, commonly known as unsupervised accuracy estimation.
@@
Current studies mainly focus on outputs or feature representation to derive a test error estimation score. Such a score can represent the calibrated test error or the distribution discrepancy between training and test datasets (Hendrycks & Gimpel, 2016; Guillory et al., 2021; Garg et al., 2022; Deng & Zheng, 2021; Yu et al., 2022b; Lu et al., 2023). For instance, Hendrycks & Gimpel (2016) considered the average maximum softmax score of the test samples as the estimated error. Similarly, Garg et al. (2022) proposed to learn a confidence threshold from the training distribution. Deng & Zheng (2021) quantified the distribution difference between training and test datasets in the feature space, while Yu et al. (2022b) gauges the distribution gap at the parameter level. Although insightful, the current body of literature overlooks a potential tool for test accuracy estimation, namely the gradients that are known to correlate strongly with the generalization error of the deep neural networks (Li et al., 2019; An et al., 2020).
@@
Recently, increasing attention has been paid to the intrinsic properties of gradients to design learning algorithms for meta-learning (Finn et al., 2017), domain generalization (Shi et al., 2021; Mansilla et al., 2021) or to improve optimization of DNNs Zhou et al. (2020); Zhao et al. (2022) by relying on them. In domain shift scenarios, Mansilla et al. (2021) proposed to clip the conflicting gradients and introduced a strategy to promote gradient agreement across multiple domains. Meanwhile, Zhao et al. (2022) designed a gradient-based regularization term to make the optimizer find flat minima. In the field of OOD detection, Huang et al. (2021) introduced a gradient-based function to detect out-of-distribution (OOD) samples (relation to our method is compared in Appendix 7). Although the works mentioned above showed the potential of gradients to tackle the learning problems both in OOD and in-distribution (ID) settings, it is still unclear how unsupervised gradient-based scores can correlate with the test accuracy and be used to estimate it. This motivates us to ask:
Are gradients predictive of test accuracy under distribution shift?
In this paper, we shed new light on this open question: we surprisingly observe that there exists a strong linear relationship between gradient norm and test accuracy under distribution shift. Moreover, our theoretical analysis shows that the gradient norm conveys information on the generalization capacity of a well-calibrated model. In a nutshell, this work provides direct evidence that gradient-based information correlates with generalization performance, which paves the way to a better understanding of how neural networks generalize across unseen domains.
@@
We hypothesize that the model requires a gradient step of large magnitude when it fails to generalize well on test data from unseen domains. To quantify the magnitude of gradients, we propose a simple yet efficient gradient-based statistic, GdScore, which employs the norm of the gradients backpropagated from a standard cross-entropy loss on the test samples. To avoid the need for ground-truth labels, we propose a pseudo-labeling strategy that benefits from both correct and incorrect predictions. We demonstrate that the norm of this one-step gradient of the classification layer strongly correlates with the generalization performance under diverse distribution shifts, acting as a strong and lightweight proxy for the latter. The main contributions of our paper are summarized as follows:
@@
    We first provide several theoretical insights showing that correct pseudo-labeling and gradient norm have a direct impact on the test error estimation. This is achieved by looking at the analytical expression of the gradient after one backpropagation step over the pre-trained model on test data under distribution shift and by upper-bounding the target out-of-distribution risk.
@@
    Based on these theoretical insights, we propose the GdScore, which gauges the magnitude of the classification-layer gradients and presents a strong correlation with test accuracy. Our method does not require access to either test labels or training datasets and only needs one step of backpropagation which makes it particularly lightweight in terms of computational efficiency compared to other self-training methods.
@@
    We demonstrate the superiority of GdScore with a large-scale empirical evaluation. We achieve new state-of-the-art results on 11 benchmarks across diverse distribution shifts compared to 8 competitors, while being faster than the previous best baseline.
Organization of the paper.
The rest of the paper is organized as follows. Section 3 presents the necessary background on the problem at hand. In Section 4, we derive the theoretical insights that motivate the GdScore introduced afterward. Section 5 is devoted to extensive empirical evaluation of our method, while the ablation study is deferred to Section 6. Finally, Section 8 concludes our work.
2 Related Work
Unsupervised accuracy estimation.
Unsupervised accuracy estimation is a vital topic in practical applications due to frequent distribution shifts and the unavailability of ground-truth labels for test samples. To comprehensively understand this field, we introduce two main existing settings related to this topic.
  1. 1.
    Some works aim to estimate the test accuracy or gauge the accuracy discrepancy between the training and the test set only via the training data (Corneanu et al., 2020; Jiang et al., 2019; Neyshabur et al., 2017; Unterthiner et al., 2020; Yak et al., 2019; Martin & Mahoney, 2020). For example, the model-architecture-based algorithm (Corneanu et al., 2020) derives plenty of persistent topology properties from the training data, which can identify when the model learns to generalize to unseen datasets. However, those algorithms are deployed under the assumption that the training and the test data are from the same distribution, which means they are vulnerable to distribution shifts.
  2. 2.
    Our work belongs to the second setting, which aims to estimate the classification accuracy of a specific test dataset during evaluation using unlabeled test samples and/or labeled training datasets. The main research direction is to explore the negative relationship between the distribution discrepancy and model performance from the space of features (Deng & Zheng, 2021), parameters (Yu et al., 2022b) and labels (Lu et al., 2023). Another popular direction is to design an estimation score via the softmax outputs of the test samples (Guillory et al., 2021; Jiang et al., 2021; Guillory et al., 2021; Garg et al., 2022), which heavily relies on model calibration. Some works also learn from the field of unsupervised learning, such as agreement across multiple classifiers (Jiang et al., 2021; Madani et al., 2004; Platanios et al., 2016; 2017) and image rotation (Deng et al., 2021). In addition, the property of the test datasets presented during evaluation has also been studied recently (Xie et al., 2023). To the best of our knowledge, our work is the first to study the linear relationship between the gradients and model performance.
Gradients in generalization. The role of gradients in generalization has attracted increasing attention recently. To gauge the generalization performance of the hypothesis learned from the training data on unseen samples, known as the out-of-sample error (Hardt et al., 2016; London, 2017; Rivasplata et al., 2018), many studies try to provide a tight upper bound for generalization error from the view of gradient descent theoretically, indicating that gradients correlate with the discrepancy between the empirical loss and the population loss (Li et al., 2019; Chatterjee, 2020; Negrea et al., 2019; An et al., 2020). However, those works assume that seen to unseen data are from the identical distribution, while unsupervised accuracy estimation discusses a more complex and realistic issue that they come from different distributions. Under distribution shift, gradients are also explored. For example, Mansilla et al. (2021) clips the conflicting gradients emerging in domain shift scenarios and promotes gradient agreement across multiple domains via the gradient agreement strategy. Similarly, Zhao et al. (2022) designs a gradient-based regularization term to make the optimizer find the flat minima. In out-of-distribution (OOD) detection which goal is to determine whether a given sample is in-distribution (ID) or out-of-distribution (Hendrycks & Gimpel, 2016; Hendrycks et al., 2018; Liu et al., 2020; Yang et al., 2021; Liang et al., 2017), (Huang et al., 2021) finds that ID data usually have higher gradient magnitude than OOD data from current source distribution to a uniform distribution. However, despite their empirical success, the relationship between gradients and generalization is still unclear.
3 Background
@@
where $\mathrm{s}_{\bm{\omega}}^{(k)}$ denotes the output of the softmax for the class $k$ approximating the posterior probability $P(Y\!=\!k|\mathbf{x})$, i.e., $\mathrm{s}_{\bm{\omega}}^{(k)}(f_{\mathbf{g}}(\mathbf{x}))=\exp\{\mathbf{w}_{k}^{\intercal}f_{\mathbf{g}}(\mathbf{x})\}/\left(\sum_{\tilde{k}}\exp\{\mathbf{w}_{\tilde{k}}^{\intercal}f_{\mathbf{g}}(\mathbf{x})\}\right)$.
@@
We now assume to have access to $m$ test samples from the target distribution $\mathcal{D}_{\text{test}}=\{\tilde{\mathbf{x}_{i}}\}_{i=1}^{m}\sim P_{T}(\mathbf{x})$, where $P_{T}(\mathbf{x},y)\neq P_{S}(\mathbf{x},y)$. For each test sample $\tilde{\mathbf{x}}_{i}$, we predict the label by $\tilde{y}^{\prime}_{i}=\operatorname*{arg\,max}_{k\in\mathcal{Y}}f_{\bm{\theta}}(\tilde{\mathbf{x}}_{i})$. We now want to assess the performance of $f_{\bm{\theta}}$ on a target distribution without using corresponding ground-truth labels $\{\tilde{y}_{i}\}_{i=1}^{m}$ by estimating as accurately as possible the following quantity:
@@
4 On the strong correlation between gradient norm and test accuracy
We start by deriving an analytical expression of the gradient obtained when fine-tuning a source pre-trained model on new test data. We further use the intuition derived from it to propose our test accuracy estimation score and justify its effectiveness through a more thorough theoretical analysis.
@@ 4 On the strong correlation between gradient norm and test accuracy
Below, we follow the setup considered by Denevi et al. (2019); Balcan et al. (2019); Arnold et al. (2021) to develop our intuition behind the importance of gradient norm in unsupervised accuracy estimation. Our main departure point for this analysis is to consider fine-tuning: a popular approach to adapting a pre-trained model to different labeled datasets is to update either all or just a fraction of its parameters using gradient descent on the new data. To this end, let us consider the following linear regression example, where the test data from unseen domains are distributed as $X\!\sim\!\mathcal{N}(0,\sigma^{x}_{t})$, $(Y|X\!=\!x)\!\sim\!\mathcal{N}(\theta_{t}x,1)$ parameterized by the optimal regressor $\theta_{t}\in\mathbb{R}$, while the data on which the model was trained is distributed as $X\!\sim\!\mathcal{N}(0,\sigma^{x}_{s})$, $(Y|X\!=\!x)\!\sim\!\mathcal{N}(\theta_{s}x,1)$ with $\theta_{s}\in\mathbb{R}$. Consider the least-square loss over the test distribution:
@@
$\frac{1}{2}\nabla_{c}\mathbb{E}_{P_{T}(x)}\mathbb{E}_{P_{S}(y|x)}[(y-cx)^{2}]$ $=\mathbb{E}_{P_{T}(x)}\mathbb{E}_{P_{S}(y|x)}[(y-cx)(-x)]$
@@
4.1 Proposed approach: GdScore
We now formally introduce our proposed score, termed GdScore, to estimate test accuracy in an unsupervised manner during evaluation. We start by recalling the backpropagation process of the pre-trained neural network $f_{\theta}$ from a cross-entropy loss and then describe how to leverage the gradient norm for the unsupervised accuracy estimation. The detailed algorithm can be found in Appendix B.
@@ 4.1 Proposed approach: GdScore
where $\tau$ denotes the threshold value, and $U[1,k]$ denotes the discrete uniform distribution with outcomes $\{1,\dots,K\}$. In a nutshell, we assign the predicted label to $\tilde{\mathbf{x}}_{i}$, when the prediction confidence is larger than a threshold while using a randomly sampled label from the label space otherwise. The detailed empirical evidence justifying this choice is shown in Section 6, and we discuss the choice of proper threshold $\tau$ in Appendix 6. From the theoretical point of view, our approach assumes that the classifier makes mistakes mostly on data with low prediction confidence, for which we deliberately assign noisy pseudo-labels. Feofanov et al. (2019) used a similar approach to derive an upper bound on the test error and proved its tightness in the case where the assumption is satisfied. We discuss this matter in more detail in Appendix E.
@@
To estimate our score, we calculate the gradients w.r.t. the weights of the classification layer $\bm{\omega}$ during the first epoch backpropagated over the standard cross-entropy loss defined by:
@@
where each unlabeled instance $\mathbf{\tilde{x}}_{i}$ is pseudo-labeled following Eq. 5. Then, the gradient of the classification layer $\bm{\omega}$ is evaluated as follows:
@@
Note that our method requires neither gradients of the whole parameter set of the pre-trained model nor iterative training. This makes it highly computationally efficient.
@@
where $||\cdot||_{p}$ denotes $L_{p}$-norm.
@@ 4.2 Theoretical analysis
In this section, we provide theoretical insights into our method. We first clarify the connection between the true target cross-entropy error and the norm of the gradients. Then, we show that the gradient norm is upper-bounded by a weighted sum of the norm of the inputs.
@@
where $\bm{\omega}=(\mathbf{w}_{k})_{k=1}^{K}\in\mathbb{R}^{D\times K}$ are the parameters of the linear classification layer $f_{\bm{\omega}}$. For the ease of notation, the gradient of $\mathcal{L}_{T}$ w.r.t $\bm{\omega}$ is denoted by $\nabla\mathcal{L}_{T}$.
The following theorem, whose proof we defer to Appendix H.1, makes the connection between the true risk and the $L_{p}$-norm of the gradient explicit.
@@
The left-hand side here is the difference in terms of the true risks obtained for the same distribution with two different classifiers. The right-hand side shows how this difference is controlled by the maximum gradient norm over the two classifiers and a term capturing how far the two are apart.
In the context of the proposed approach, we want to know the true risk of the source classifier $\bm{\omega}_{s}$ on the test data and its change after one step of gradient descent. The following corollary applies Theorem 4.1 to characterize this exact case. The proof is deferred to Appendix H.2.
@@
$|\mathcal{L}_{T}(\bm{\omega}_{s})\!-\!\mathcal{L}_{T}(\mathbf{c})|$ $\!\leq\!\eta\max_{\bm{\omega}}(\left\lVert\nabla\mathcal{L}_{T}(\bm{\omega})\right\rVert_{p})\left\lVert\nabla\mathcal{L}_{T}(\bm{\omega}_{s})\right\rVert_{q}.$
Note that in this case $\mathcal{L}_{T}(\bm{\omega}_{s})$ can be seen as a term providing the true test risk after pseudo-labeling it with the source classifier. This shows the importance of pseudo-labeling as it acts as a departure point for obtaining a meaningful estimate of the right-hand side. When the latter is meaningful, the gradient norm on the right-hand side controls it together with a magnitude that tells us how far we went after one step of backpropagation.
In the next theorem, we provide an upper bound on the $L_{p}$-norm of the gradient as a weighted sum of the $L_{p}$-norm of the inputs. The proof is deferred to Appendix H.3.
@@
For any $p\geq 1$ and $\forall\,\bm{\omega}\in\mathbb{R}^{D\times K}$, the $L_{p}$-norm of the gradient can be upper-bounded as follows:
$\left\lVert\nabla\mathcal{L}_{T}(\bm{\omega})\right\rVert_{p}$ $\leq\mathbb{E}_{P_{T}(\mathbf{x},y)}\alpha(\bm{\omega},\mathbf{x},y)\cdot\left\lVert\mathbf{x}\right\rVert_{p},$
@@
Hence, the norm of the gradient is upper-bounded by a weighted combination of the norm of the inputs, where the weight $\alpha(\bm{\omega},\mathbf{x},y)\in[0,1]$ conveys how well the model predicts on $\mathbf{x}$. In the case of perfect classification, the upper bound is tight and equals 00. In practice, as we do not have access to the true risk, the gradients can be approximated by the proposed GdScore that requires to pseudo-label test data by Eq. 5. As we said earlier, this implies that the model has to be well calibrated (see Appendix E), which is conventional to assume for self-training methods (Amini et al., 2023) including the approach of Yu et al. (2022b). Then, the network projects test data into the low confidence regions, and the gradient for these examples will be large as we need to update $\bm{\omega}_{s}$ significantly to fit them.